Jose Reynolds deposited $10,000 in an account paying interest of 4% compounded annually Click here to view the factor table FWingate Company borrowed $9O,000 on January 2, 2019. This amount plus accrued interest of 6% compounded annually will be repaClick here to view the factor table Future Value of 1 Click here to view the factor table Future Value of an Annuity of 1 CliFrye Company is considering investing in an annuity contract that will return $50,000 annually at the end of each year for 20TABLE 1Future Value of 1 (n) Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% 0 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000 1.00TABLE 2 Future Value of an Annuity of 1 (n) Payments 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% 1.00000 1.0000 1.00000 1 1.00000 1.000TABLE 3 Present Value of 1 (n) Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% 1 96154 95238 9434093458 92593 91743 90909 90090 892Present Value of an Annuity of 1 TABLE 4 (п) |Pаyments 4% 5% 6% 7% 8% 9% 10% 11% 12% 15% 1 96154 95238 94340 93458 92593 9174
Answer #1

Ans 1:

Principal, P = $10,000

Rate of interest, r = 4% compounded annually

Time period, n = 4

Amount, A = P (1 + r)n = $10,000 (1 + 0.04)4 = $10000 x 1.16986 (1.16986 is the value for r = 4% & n=4 in table 1)

So, Amount in the account at the end of four years = $11698.60

Ans 2:

Principal, P = $90,000

Rate of interest, r = 6% compounded annually

Time period, n = 3

Amount, A = P (1 + r)n = $90,000 (1 + 0.06)3 = $90000 x 1.19102 (1.19102 is the value for r = 6% & n=3 in table 1)

So, Amount Wingate repay at the end of third year = $11910.20

Ans 3:

Future Value, F.V. = $90,000

Discount rate, d = 9%

Time periods, n = 7

Present Value, P.V. = F.V. / (1+d)n = $90000 / (1+0.09)7

= $90000 x 0.54703 (0.54703 is the discounting factor for r = 9% and n=7 from table 3)

= $49232.70

So, Present value = $49232.70

Ans 4:

Future Value, F.V. = $150,000

Discount rate, d = 12%

Time periods, n = 5

Present Value, P.V. = F.V. / (1+d)n = $150000 / (1+0.12)5

= $150000 x 0.56743 (0.56743 is the discounting factor for r = 12% and n=5 from table 3)

= $85114.50

So, Present value = $85114.50

Ans 5:

Annuity Amount, A = $50000

Annuity period, n = 20

Rate of return, r = 8%

Present Value of annuity, P.V. = A x [ ((1-(1+r)-n) / r ]

= A x (discount factor for r=8% & n=20 in table 4)

=$50000 x 9.81815

=$490907.37

So, the amount Frye company pay for investment = $490907.37